# Symbolic Learning of Interpretable Reduced-Order Models for Jumping Quadruped Robots

### v1

@@ Abstract
Reduced-order models are essential for motion planning and control of quadruped robots, as they simplify complex dynamics while preserving critical behaviors. This paper introduces a novel methodology for deriving such interpretable dynamic models, specifically for jumping. We capture the high-dimensional, nonlinear jumping dynamics in a low-dimensional latent space by proposing a learning architecture combining Sparse Identification of Nonlinear Dynamics (SINDy) with physical structural priors on the jump dynamics. Our approach demonstrates superior accuracy to the traditional actuated Spring-loaded Inverted Pendulum (aSLIP) model and is validated through simulation and hardware experiments across different jumping strategies.
@@
Quadruped Robots, Dynamic Modeling, Symbolic Regression, Autoencoders, Interpretability, Reduced-order Models.
@@ 1 Introduction
Recently, quadruped robots have garnered significant attention due to their versatility and potential applications in various fields, including search and rescue missions, agricultural tasks, and exploration in hazardous environments (Gong et al., 2023; Sombolestan and Nguyen, 2024). The performance of these robots depends on accurate dynamic modeling to capture their complex locomotion behaviors.
Traditional modeling techniques, such as the Lagrangian and Newton-Euler methods, have been widely used to derive high-fidelity whole-body dynamic models for quadruped robots (Bellicoso et al., 2017; Di Carlo et al., 2018; Yan et al., 2021; Ding et al., 2024b). However, these high-dimensional models are often computationally intractable for real-time applications due to their complexity (Li et al., 2023). To address this, Researchers have developed reduced-order template models that approximate essential dynamics while significantly lowering computational costs. Examples include the famous Spring-Loaded Inverted Pendulum (SLIP) model (Blickhan, 1989) and the Single Rigid Body (SRB) (Hong et al., 2020), which simplifies multi-body interactions into low-dimensional representations, enabling efficient analysis of gaits like running and hopping. However, such simplifications often sacrifice accuracy, particularly in scenarios involving multi-contact transitions, nonlinear limb dynamics, or complex terrain interactions (Spröwitz et al., 2013; Carpentier and Mansard, 2018).
To tackle this issue, we investigate the possibility of using a data-driven strategy to learn reduced-order models directly, with a particular interest in the compact, interpretable, and precise description of jumping dynamics, which is very challenging. Numerous studies have utilized machine learning and deep learning methods to model the complex dynamics of various systems (Calandra et al., 2015; Ogunmolu et al., 2016; Rajendra and Brahmajirao, 2020; Hashemi et al., 2023). Dimensionality reduction techniques like autoencodershave been employed to extract low-dimensional features in a latent space (Hinton and Salakhutdinov, 2006; Gonzalez and Balajewicz, 2018; Refinetti and Goldt, 2022), and then Multilayer Perceptrons (MLPs) or Recurrent Neural Networks (RNNs) can learn the dynamics in this low-dimensional space, but at the expense of interpretability in exchange for accuracy. Recent hybrid approaches that combine physics with learning (Stölzle, 2025), such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), Deep Lagrangian Networks (DeLaNs) (Lutter et al., 2019; Liu et al., 2024), Coupled Oscillator Networks (CONs) (Stölzle and Santina, 2024), integrate physical laws with deep learning, enhancing accuracy, interpretability, data efficiency, and out-of-distribution performance. However, PINNs often require physics knowledge that is only available to domain experts. Furthermore, nonlinear time-dependent behaviors, as for example experienced during contact phases with the environment, are usually not included in the physical priors and can, therefore, not be learned by the DeLaNs. Symbolic regression, particularly SINDy (Brunton et al., 2016; Pandey et al., 2024), presents an alternative through interpretable symbolic equations, though their application to quadruped robots remains unexplored.
This work proposes a novel methodology to derive reduced-order, interpretable dynamic models for jumping robots. Building upon SINDy autoencoders (Champion et al., 2019), our approach uses a linear  [^1] The linearity of the autoencoder is critical for preserving interpretability and facilitating the closed-form, analytical (i.e., without autodifferentiation) computation of time derivatives. autoencoder with shared weights to reduce the dimension of the state space while keeping configurations and maintaining their derivatives distinct in the latent space. We map inputs to latent space through the trained autoencoder, optimizing its weights concurrently with the SINDy latent model using an ad-hoc loss function. Considering the hybrid nature of the system during different jumping phases, we learn a consistent model across various hybrid phases by first training the encoder in one phase (e.g., the launching and landing phase), then optimizing the decoder across all phases, and finally learning the SINDy parameters with frozen autoencoder weights.
The main contributions of this work are threefold:
  • We present a novel modeling approach that integrates SINDy with a linear autoencoder, utilizing shared weights to derive reduced-order, interpretable models for quadruped robots. This method enhances computational efficiency and model interpretability.
  • We propose a sequential training strategy to accurately model the distinct jump phases, ensuring a consistent model across various hybrid phases.
  • We evaluate our novel approach on various quadruped robots with different jumping modalities to demonstrate versatility.
The rest of this paper follows: Section 2 covers preliminary concepts; Section 3.1 discusses the proposed learning architecture; Section 3 describes the remaining aspects of our experimental methodology; Section 4 evaluates the method, featuring simulation results for various quadruped models, latent dimensionality analysis, and hardware validation. Section 5 concludes this work.
@@ 2 Preliminary
Traditional modeling approaches, such as multi-rigid-body dynamics, represent quadruped robots as a movable base (i.e., the body) with kinematic chains (i.e., the legs) attached, applicable to various quadruped robots, as shown in Fig. 1.
@@
As illustrated in Fig. 1, the Unitree Go1, for example, exhibits 121212 Degrees of Freedom (DOFs) for the limbs alone, with 666 underactuated DoFs for the floating base’s position and orientation relative to an inertial frame (Righetti et al., 2011).
The configuration space of quadruped robots is typically parameterized as (Mistry et al., 2010)
@@ 3.1 Proposed Learning Architecture
We present a novel framework combining linear autoencoders with SINDy to model quadruped dynamics, incorporating physical priors for dimensionality reduction while preserving essential dynamic features. Our approach achieves dimensionality reduction while preserving the essential dynamic characteristics, enabling efficient representation of complex jumping motions
Fig. 2 provides an overview of the encoded latent space and the established dynamics loop.
Figure 2: Overview of our proposed methodology: We learn symbolic reduced-order dynamics $\ddot{\bm{\xi}}=\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi}$ that operate in a low-dimensional space $(\bm{\xi},\dot{\bm{\xi}})$. This allows us to predict the time evolution of the higher-dimensional system states $(\mathbf{q},\dot{\mathbf{q}})$. Here, $\mathbf{q}_{0}$ and $\dot{\mathbf{q}}_{0}$ represent the initial state.
Fig. 3 details the training process, including the encoder-decoder structure, latent space dynamics prediction, and real-space dynamics reconstruction. The algorithm includes encoding the initial state into a latent space, predicting dynamics in this latent space, and decoding them back to the full-order system space. We ensure accurate reconstruction of system states and dynamical modeling through specialized loss components.
@@
Figure 3: Detailed visualization of the training methodology. (a): The linear autoencoder is first trained on the contact phase dataset, minimizing the reconstruction loss, $\mathcal{L}_{\mathrm{rec}}$, depicted by the black line. Subsequently, in the second step (visualized with a blue line in the flowchart), the encoder weights are kept fixed, and the decoder is fine-tuned on a dataset that includes all motion phases. (b): In the third step, we train, for each motion phase separately, a symbolic dynamical model $\bm{\ddot{\xi}}_{\text{pred}}$ in latent space. Here, we employ the SINDy approach to regress a sparse matrix $\bm{\Xi}$ that selects and scales basis functions from the library $\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The loss components $\mathcal{L}_{\text{SINDy}}$ and $\mathcal{L}_{\text{reg}}$ ensure accurate dynamics prediction and sparsity of the coefficients, respectively. The SINDy loss $\mathcal{L}_{\text{SINDy}}$ is calculated based on the predicted reduced-order model accelerations, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, and the corresponding decoded configuration-space accelerations, $\ddot{\mathbf{q}}_{\mathrm{pred}}$. Importantly, we keep the autoencoder weights frozen, which preserves training stability and ensures that the reconstructions remain accurate during all motion phases.
The linear autoencoder is designed to reduce the dimensionality of the state space while preserving a clear linear relationship with the original space, thereby maintaining direct interpretability. The autoencoder consists of an encoder $\left(\mathbf{W}_{\mathrm{e}},\mathbf{B}_{\mathrm{e}}\right)$ and a decoder $\left(\mathbf{W}_{\mathrm{d}},\mathbf{B}_{\mathrm{d}}\right)$. The encoder’s weights and bias are $\mathbf{W}_{\mathrm{e}}\in\mathbb{R}^{l\times(m+6)}$ and $\mathbf{B}_{\mathrm{e}}\in\mathbb{R}^{l}$, while the decoder’s weights and bias are $\mathbf{W}_{\mathrm{d}}\in\mathbb{R}^{(m+6)\times l}$ and $\mathbf{B}_{\mathrm{d}}\in\mathbb{R}^{m+6}$, where $l$ is the dimension of the desired latent space configuration.
The encoder maps the robot’s configuration $\mathbf{q}$ and its derivatives $\mathbf{q}^{(n)}$ into a lower-dimensional latent space using the linear transformation
@@
where $\bm{\xi}\in\mathbb{R}^{l}$ represents the latent representation, $\bm{\xi}^{(n)}$ denotes the $n$-th derivative of $\bm{\xi}$, and $\delta_{n0}$ is the Kronecker delta, equal to 111 when $n=0$ and 0 otherwise. The system’s actuation is mapped into latent space as
@@
After establishing the latent space mapping, we employ the SINDy algorithm (Brunton et al., 2016) to identify symbolic representations of the latent dynamics (see Fig. 3(b)). SINDy constructs a symbolic library of candidate functions $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$, including polynomial and trigonometric functions of latent variables ($\bm{\xi}$), their derivatives ($\bm{\xi}^{(}n)$), and the transformed control inputs ($\bm{\nu}$). This library represents a comprehensive set of terms that might appear in the system’s governing equations. The objective is to represent the dynamics of the latent variables as a sparse combination of these candidate functions. Mathematically, this is expressed as a sparse regression problem, with the latent dynamics expressed as
where $\bm{\Xi}\in\mathbb{R}^{p\times l}$ is a sparse matrix of coefficients to be determined, with $p$ being the number of the candidate functions in $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The optimization problem is typically solved using sequential thresholding methods to ensure a sparse solution (Brunton et al., 2016; Champion et al., 2019). Promotes sparsity by penalizing the $\ell_{1}$-norm of $\bm{\Xi}$, which encourages most entries in $\Xi$ to be zero. This sparsity-promoting optimization ensures that only a minimal set of candidate functions is selected, resulting in an interpretable and computationally efficient model.
Subsequently, the predicted accelerations in the latent space, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, are transformed back to the original state space by the linear decoder
@@
For training encoder, decoder, and reduced-order symbolic dynamics, we consider the following loss components:
@@
3.2 Training with Multi-Phase Motions
Jumping motion poses unique modeling challenges due to its hybrid dynamical nature. This study considers two types of jumping: the synchronized and the froggy jumps.
In the synchronized jump, the robot launches and lands with full limb contact simultaneously. This motion consists of two dominant phases: (1) a contact phase, where all legs remain grounded during takeoff and landing, and (2) a flight phase, where the robot is entirely airborne.
Conversely, the froggy jump consists of an initial phase where all legs are grounded, followed by a transition phase where only the rear legs remain in contact with the ground, and then progressing to a flight phase.
To effectively model the hybrid jumping dynamics, we propose a structured training methodology (Fig. 3) that combines transfer learning principles (Torrey and Shavlik, 2010) with phase-specific optimization. First, we train the encoder and decoder statically on reconstructing the configurations of the full contact phase by minimizing the $\mathcal{L}_{\mathrm{recon}}$ loss. Then, the encoder weights are frozen, and the decoder is trained on all motion phases, including the flight phase, still minimizing $\mathcal{L}_{\mathrm{recon}}$. Finally, with the encoder and decoder weights fixed, we train the symbolic dynamics for each motion phase (contact, partial contact, and flight) by minimizing the $\mathcal{L}_{\mathrm{SINDy}}$ and $\mathcal{L}_{\mathrm{reg}}$ loss terms, efficiently capturing the hybrid nature of the jumping dynamics.
This training methodology ensures that the final model is physically coherent and can accurately reconstruct the robot’s motion across different phases. By leveraging transfer learning, we create robust models that effectively capture hybrid dynamics, enhance interpretability, and seamlessly handle complex, phase-dependent behaviors.

### v2

@@ Abstract
Reduced-order models are central to motion planning and control of quadruped robots, yet existing templates are often hand-crafted for a specific locomotion modality. This motivates the need for automatic methods that extract task-specific, interpretable low-dimensional dynamics directly from data. We propose a methodology that combines a linear autoencoder with symbolic regression to derive such models. The linear autoencoder provides a consistent latent embedding for configurations, velocities, accelerations, and inputs, enabling the sparse identification of nonlinear dynamics (SINDy) to operate in a compact, physics-aligned space. A multi-phase, hybrid-aware training scheme ensures coherent latent coordinates across contact transitions. We focus our validation on quadruped jumping—a representative, challenging, yet contained scenario in which a principled template model is especially valuable. The resulting symbolic dynamics outperform the state-of-the-art handcrafted actuated spring-loaded inverted pendulum (aSLIP) baseline in simulation and hardware across multiple robots and jumping modalities.
@@
Quadruped jumping, Symbolic Regression, Autoencoders, Interpretability, Reduced-order Model.
@@ 1 Introduction
Legged robots are now pervasive systems, deployed in applications ranging from planetary exploration (Arm et al., 2023) to search-and-rescue operations (Bai et al., 2018). Yet, despite this growing presence, their dynamics remain remarkably difficult to model—especially when seeking a description expressive enough to capture essential behaviors and lightweight enough for real-time use.
Classic Lagrangian and Newton–Euler formulations provide full-order models of legged systems (Bellicoso et al., 2017; Di Carlo et al., 2018; Yan et al., 2021; Ding et al., 2024b), but these models remain too computationally heavy for online deployment (Li et al., 2023). This has led to reduced-order templates such as spring-loaded inverted pendulum (SLIP) (Blickhan, 1989) and single rigid body dynamics (SRBD) (Hong et al., 2020), which capture dominant behaviors at far lower cost. Yet these templates lose accuracy in multi-contact transitions, rapid reconfiguration, and complex terrain (Carpentier and Mansard, 2018). In practice, one must choose between high-dimensional models that are difficult to deploy and compact, task-specific heuristics. A principled way to automatically extract reduced-order models—accurate, compact, and interpretable—would significantly advance the analysis and control of hybrid locomotion dynamics. However, such a method is still missing.
In this paper, we turn to interpretable data-driven techniques to address this challenge. Data-driven learning has shown strong capabilities in modeling physical systems, with numerous works leveraging neural networks to capture complex dynamics (Calandra et al., 2015; Ogunmolu et al., 2016; Rajendra and Brahmajirao, 2020; Hashemi et al., 2023). Dimensionality-reduction methods such as autoencoders extract low-dimensional latent features (Hinton and Salakhutdinov, 2006; Gonzalez and Balajewicz, 2018; Refinetti and Goldt, 2022), after which multi-layer perceptrons (MLPs) or recurrent neural networks (RNNs) can model reduced-order dynamics—often gaining accuracy at the expense of interpretability. Physics-informed approaches, including physics-informed neural networks (PINNs) (Raissi et al., 2019), deep Lagrangian networks (DeLaN) (Lutter et al., 2019; Liu et al., 2024), and coupled oscillator networks (Stölzle and Santina, 2024), improve interpretability, data efficiency, and out-of-distribution performance. Still, none of these methods handle the hybrid, time-dependent structure inherent to locomotion. This limitation is particularly acute for floating-base dynamics, where contacts induce abrupt changes in inertial and potential forces—effects that current learning frameworks are not designed to capture.
Given this gap, no existing work provides a systematic method for learning reduced-order models of legged systems. We address this by combining linear autoencoders with symbolic regression, specifically the sparse identification of nonlinear dynamics (SINDy) (Brunton et al., 2016; Pandey et al., 2024), extending prior autoencoder-based formulations (Champion et al., 2019) to quadruped dynamics, where such techniques remain unexplored.
The linear autoencoder is essential: its constant Jacobian—equal to its weight matrix—provides a unified linear mapping of configurations, velocities, accelerations, and inputs, enabling consistent and interpretable identification of reduced-order dynamics. Direct identification in the observation space is impractical due to scale disparities and dimensionality, while nonlinear autoencoders introduce state-dependent latent geometries and distort physical structure through separate nonlinear mappings of positions and velocities.
To obtain a representation valid across hybrid phases, we adopt a multi-phase training scheme: train the encoder in the contact phase (launch + landing), refine the decoder across all phases, and finally identify SINDy parameters with the autoencoder frozen. We validate the approach on jumping motions, a demanding form of dynamic locomotion that enables navigation in complex environments (Nguyen et al., 2019; Atanassov et al., 2024; Ding et al., 2023) yet remains challenging due to high accelerations and impact-driven dynamics.
In summary, this work contributes to legged robotics and physics-informed machine learning through:
  • A physics-aligned linear latent embedding that provides consistent linear mappings of positions, velocities, accelerations, and inputs, enabling symbolic learning of reduced-order and interpretable jumping dynamics.
  • A sequential multi-phase training strategy that preserves coherent latent coordinates and symbolic models across the hybrid phases of a jump.
  • The first demonstration of learning reduced-order dynamics for legged locomotion, with evaluation across different quadruped robots and jumping modalities.
@@ 2 Preliminary
2.1 Robot description
Traditional modeling approaches represent the quadruped robot as a movable base (i.e., the body) with kinematic chains (i.e., the legs) attached, as shown in Fig. 1.
@@ 2.1 Robot description
As illustrated in Fig. 1, the Unitree Go1, for example, exhibits $12$ degrees of freedom (DOFs) for the limbs alone, with $6$ underactuated DoFs for the floating base’s position and orientation relative to an inertial frame (Righetti et al., 2011). The configuration space of quadruped robots is typically parameterized as (Mistry et al., 2010)
@@
2.2 Encoder–Decoder Representation Learning
The autoencoder, i.e., encoder–decoder architecture, enables one to obtain a compact but informative representation of high-dimensional state variables (Hinton and Salakhutdinov, 2006). The encoder $E(\cdot)$ maps the original variable $\bm{x}\in\mathbb{R}^{n}$ into a low-dimensional latent vector $\bm{z}=E(\bm{x})\in\mathbb{R}^{k}$ with $k\ll n$, effectively capturing the essential features that characterize the system’s dynamics. The decoder $D(\cdot)$ reconstructs the high-dimensional signal from the latent representation, i.e., $\hat{\bm{x}}=D(\bm{z})$, ensuring that the latent space preserves the underlying structure of the original data. This framework not only reduces computational complexity but also provides a smooth and continuous latent manifold that is amenable to downstream tasks such as policy learning.
2.3 SINDy
As a typical symbolic learning approach, SINDy (Brunton et al., 2016) enables us to identify symbolic representations of the dynamics by constructing a symbolic library of candidate functions $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$, including polynomial and trigonometric functions of latent variables ($\bm{\xi}$), their derivatives ($\bm{\xi}^{(n)}$), and the transformed control inputs ($\bm{\nu}$). This library represents a comprehensive set of terms that might appear in the system’s governing equations. The objective is to represent the dynamics of the latent variables as a sparse combination of these candidate functions. Mathematically, this is expressed as a sparse regression problem, with the latent dynamics expressed as
where $\bm{\Xi}\in\mathbb{R}^{p\times l}$ is a sparse matrix of coefficients to be determined, with $p$ being the number of the candidate functions in $\Theta(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The optimization problem is typically solved using sequential thresholding methods to ensure a sparse solution (Brunton et al., 2016; Champion et al., 2019). Promotes sparsity by penalizing the $\ell_{1}$-norm of $\bm{\Xi}$, which encourages most entries in $\Xi$ to be zero. This sparsity-promoting optimization ensures that only a minimal set of candidate functions is selected, resulting in an interpretable model.
@@ 3.1 Proposed Learning Architecture
We combine linear autoencoders with SINDy to model jumping dynamics in the latent space, achieving dimensionality reduction while preserving the essential dynamic characteristics, enabling efficient representation of complex jumping motions. Fig. 2 provides an overview of the encoded latent space and the established dynamics loop.
Figure 2: Methodology overview: Symbolic latent-space dynamics $\ddot{\bm{\xi}}=\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})\,\bm{\Xi}$ are learned to predict the evolution of the full-state system from $(\mathbf{q},\dot{\mathbf{q}})$.
Fig. 3 details the core components, including the encoder-decoder structure, latent space dynamics prediction, and real-space dynamics reconstruction. That is, we encode the initial state into a latent space, predict the dynamics in this latent space, and decode them back to the full-order system space. In particular, we ensure accurate reconstruction of system states and dynamical modeling through specialized loss components.
@@
Figure 3: Flowchart of the pipeline. (a) The linear autoencoder is first trained on the contact phase dataset, minimizing the reconstruction loss, $\mathcal{L}_{\mathrm{rec}}$, depicted by the black line. Subsequently, by fixing the encoder, the decoder is fine-tuned on a dataset that includes all motion phases (see the the blue line). (b) In the third step, we train, for each motion phase separately, a symbolic dynamical model $\bm{\ddot{\xi}}_{\text{pred}}$ in latent space. Here, we employ the SINDy approach to regress a sparse matrix $\bm{\Xi}$ that selects and scales basis functions from the library $\bm{\Theta}(\bm{\xi},\dot{\bm{\xi}},\bm{\nu})$. The loss components $\mathcal{L}_{\text{SINDy}}$ and $\mathcal{L}_{\text{reg}}$ ensure accurate dynamics prediction and sparsity of the coefficients, respectively. The SINDy loss $\mathcal{L}_{\text{SINDy}}$ is calculated based on the predicted reduced-order model accelerations, $\ddot{\bm{\xi}}_{\mathrm{pred}}$, and the corresponding decoded configuration-space accelerations, $\ddot{\mathbf{q}}_{\mathrm{pred}}$.
To maintain direct interpretability, we adopt the linear autoencoder to reduce the dimensionality of the state space while preserving a linear relationship with the original space. The autoencoder consists of an encoder $\left(\mathbf{W}_{\mathrm{e}},\mathbf{B}_{\mathrm{e}}\right)$ and a decoder $\left(\mathbf{W}_{\mathrm{d}},\mathbf{B}_{\mathrm{d}}\right)$. The encoder’s weights and bias are $\mathbf{W}_{\mathrm{e}}\in\mathbb{R}^{l\times(m+6)}$ and $\mathbf{B}_{\mathrm{e}}\in\mathbb{R}^{l}$, while the decoder’s weights and bias are $\mathbf{W}_{\mathrm{d}}\in\mathbb{R}^{(m+6)\times l}$ and $\mathbf{B}_{\mathrm{d}}\in\mathbb{R}^{m+6}$, where $l$ is the dimension of the desired latent space.
With this encoder, we map the robot’s configuration $\mathbf{q}$ and its derivatives $\mathbf{q}^{(n)}$ into a lower-dimensional latent space using the linear transformation
@@
where $\bm{\xi}\in\mathbb{R}^{l}$ represents the latent representation, $\bm{\xi}^{(n)}$ denotes the $n$-th derivative of $\bm{\xi}$, and $\delta_{n0}$ is the Kronecker delta, equal to $1$ when $n=0$ and 0 otherwise. The system’s actuation is mapped into latent space as
@@
After establishing the latent space mapping, we employ the SINDy algorithm to identify symbolic representations of the latent dynamics (see Fig. 3(b)). Particularly, we follow Eq. (7) to predict the accelerations. Subsequently, the predicted accelerations $\ddot{\bm{\xi}}_{\mathrm{pred}}$ are transformed back to the original state space by the linear decoder
@@
For training the encoder, decoder, and reduced-order symbolic dynamics, we consider the following loss components:
@@
3.2 Multi-Phase Training Strategy
Jumping motion poses unique modeling challenges due to its hybrid dynamical nature. This study considers two types of jumping: the pronking and the froggy jumps.
In the pronking jump, the robot launches and lands with full limb contact simultaneously. This motion consists of two dominant phases: (1) a contact phase, where all legs remain grounded during takeoff and landing, and (2) a flight phase, where the robot is entirely airborne. Conversely, the froggy jump consists of an initial phase where all legs are grounded, followed by a transition phase where only the rear legs remain in contact with the ground, and then progressing to a flight phase.
To effectively model the hybrid dynamics, we propose a structured training methodology (Fig. 3) that combines transfer learning principles (Torrey and Shavlik, 2010) with phase-specific optimization. First, we train the encoder and decoder statically on reconstructing the configurations of the full contact phase by minimizing the $\mathcal{L}_{\mathrm{recon}}$ loss. Then, the encoder weights are frozen, and the decoder is trained on all motion phases, including the flight phase, while still minimizing $\mathcal{L}_{\mathrm{recon}}$. Finally, with the encoder and decoder weights fixed, we train the symbolic dynamics for each motion phase (contact, partial contact, and flight) by minimizing the $\mathcal{L}_{\mathrm{SINDy}}$ and $\mathcal{L}_{\mathrm{reg}}$ loss terms. In this way, the final model is physically coherent and can accurately reconstruct the robot’s motion across different phases.